A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 111.5-cm and a standard deviation of 2.4-cm.
Find the probability that the length of a randomly selected steel rod is between 105.3-cm and 115.6-cm.
P(105.3-cm < X < 115.6-cm) =

Answer by stanbon(75887)

A company produces steel rods. The lengths of the steel rods are normally distributed with a mean of 111.5-cm and a standard deviation of 2.4-cm.
Find the probability that the length of a randomly selected steel rod is between 105.3-cm and 115.6-cm.
P(105.3-cm < X < 115.6-cm) =
-------
z(105.3) = (105.3-111.5)/2.4 = -2.58333
z(115.6) = (115.6-111.5)/2.4 = 1.70833
----------------
P(105.3-cm < X < 115.6-cm) = = P(-2.58333 < z < 1.70833) = 0.9513
==================
Cheers,
Stan H.
===================